OpenStudy (mikezack123):

After a discount of 10%, the price of a pair of jeans, including tax, was $69.62. The following equation can be used to find x, the price of the pair of jeans without the discount. x - 0.1x + 0.07(x - 0.1x) = 69.62 What is the price of the pair of jeans without the discount?

OpenStudy (thesecret20111):

Try working out the 0.07(x-0.1x) part first

OpenStudy (mikezack123):

0.07x-0.007

OpenStudy (mikezack123):

?

OpenStudy (thesecret20111):

Yes very good :)

OpenStudy (mikezack123):

thats all?

OpenStudy (thesecret20111):

Ok now we have x - 0.1x + 0.07x-0.007= 69.62

OpenStudy (thesecret20111):

Can you simplify that further?

OpenStudy (mikezack123):

no?

OpenStudy (thesecret20111):

x - 0.1x = 0.9x

OpenStudy (thesecret20111):

Now we have 0.9x + 0.07x-0.007= 69.62

OpenStudy (thesecret20111):

0.9x + 0.07x = 0.97x

OpenStudy (thesecret20111):

So now we have 0.97x-0.007= 69.62 now can you see what we need to do next?

OpenStudy (mikezack123):

x=0.007?

OpenStudy (thesecret20111):

Not quite.. Lets do this step by step so we don'[t miss anything

OpenStudy (mikezack123):

ok

OpenStudy (thesecret20111):

We need add 0.007 to both sides. That gives us 0.97x= 69.627

OpenStudy (mikezack123):

wait the anser is $70.58?

OpenStudy (thesecret20111):

It will be close to that, yes. I'm not sure if that'll be the exact answer though.

OpenStudy (mikezack123):

oh okay, lets do it step by step like u said. :)

OpenStudy (thesecret20111):

Ok now we need to divide both sides by 0.97 so that we get x by itself.

OpenStudy (thesecret20111):

So that would be 0.97x/0.97= 69.627/0.97 This will give us x=69.627/0.97

OpenStudy (mikezack123):

ok so its x=71.78

OpenStudy (mikezack123):

?

OpenStudy (thesecret20111):

Yes :)

OpenStudy (thesecret20111):

So now you answer in a sentence

OpenStudy (thesecret20111):

The total price of the pair of jeans is $71.78 before the discount.

OpenStudy (mikezack123):

thats odd, :/ these are the multiple choice answers $70.58 $72.29 $74.86 $80.02

OpenStudy (thesecret20111):

Sorry! I must've made an error somewhere :(

OpenStudy (thesecret20111):

I'll try to find it.

OpenStudy (mikezack123):

thanks so much! you're a big help! :)

OpenStudy (thesecret20111):

Ok so the first step is x - 0.1x + 0.07(x - 0.1x) = 69.62 = x - 0.1x + 0.063x = 69.62 next step: x - 0.1x + 0.063x = 69.62 = 0.9x+ 0.063x = 69.62 next step: 0.9x+ 0.063x = 69.62 = 0.963x=69.62 next step: 0.963x/0.963=69.62/0.963 =x = 72.29

OpenStudy (mikezack123):

Thanks so much! :) you were a big help!

OpenStudy (thesecret20111):

I've done it a different way so there must have been an error somewhere in the old working out.

OpenStudy (thesecret20111):

No problem

OpenStudy (thesecret20111):

Thanks, and good luck :)